### v1

11institutetext: Institute for Astronomy, School of Physics, Zhejiang University, Hangzhou 310027, China
11email: [llh_astro; bbliu]@zju.edu.cn
22institutetext: Department of Physics, National University of Singapore, Singapore 117542, Singapore
22email: zekai77@u.nus.edu
Resonance Capture and Stability Analysis for Planet Pairs under Type I Disk Migration
Linghong Lin 11    Beibei Liu 11⋆⋆    Zekai Zheng 1122
We present a theoretical model to investigate a two-planet pair that undergoes convergent type I migration in a gaseous protoplanetary disk and traps into the first-order mean motion resonance. Our study identifies the conditions for resonant capture and explores the subsequent dynamical stability of the system. We derive the analytical criteria for planets with an arbitrary mass ratio and validate through numerical N-body simulations. Slow migration and weak eccentricity damping are required for resonant capture, the latter of which has not received enough attention in the literature. Once capture into resonance, their following stability can be classified into three regimes: stable trap, overstable trap and escape. Notably, resonant capture remains stable when the inner planet significantly outweighs the outer one. However, the subsequent evolution can be diverse when the mass of the inner planet is lower or comparable to that of the outer planet. Stability weakens as the relative strength between migration and eccentricity damping increases. The key results can be comprehensively demonstrated in a $\tau_{a}$-$\tau_{a}/\tau_{\rm e}$ plot, where $\tau_{a}$ and $\tau_{\rm e}$ are orbital decay and eccentricity damping timescales, respectively.
Key Words.:
planets and satellites: dynamical evolution and stability – celestial mechanics – methods: analytical
1 Introduction
Low-mass planets embedded in gaseous protoplanetary disks interact with disk gas and undergo type I orbital migration (Kley & Nelson, 2012; Paardekooper et al., 2023). Resonance capture can be a natural outcome of the planet pairs featured with convergent migration (Lee & Peale, 2002). The key question is to understand the under which conditions this capture can occur. The classical pendulum model pointed out that the trapping holds when the time planet migrates across the resonance width is longer than the time resonant angle librates (Murray & Dermott, 1999; Ogihara & Kobayashi, 2013; Batygin, 2015; Liu et al., 2015). A new criterion was derived by Huang & Ormel (2023), where they analyzed the resonant equilibrium state under the disk migration effects in a restrict three-body framework. Batygin & Petit (2023) provided an equivalent criterion using the Hamiltonian theory and extended it to the context of the unrestricted problem.
On the other hand, Goldreich & Schlichting (2014) first investigated the long-term stability of the resonant pair after capture in a gaseous disk. They found that under certain conditions, the resonant equilibrium state can become overstable, indicating that the planets may temporarily or permanently escape from resonance. Deck & Batygin (2015) conducted a similar analysis based on the Hamiltonian theory. Xu & Lai (2017) extended the framework to second-order mean motion resonances. Their study highlighted that the mass ratio of the planet pair is a crucial parameter in determining the overstability. Nesvorný et al. (2022) applied the overstable libration to account for the high eccentricity and large amplitude of the resonant angle observed in TOI-216b. Recently, Afkanpour et al. (2024) verified this overstable resonant trap using hydrodynamical simulations. They found that the outcome is related to the disk properties, such as disk mass and viscosity.
In this paper, we develop a unified framework for a two-planet pair that undergoes convergent type I migration and gets trapped at the first-order mean motion resonance ($j$:$j-1$ MMR). We investigate the conditions under which planets can be captured in resonances and examine the following stability properties. The key analytical criteria are derived for planets with arbitrary mass ratios and the outcomes are summarized in a $\tau_{\rm a}{-}\tau_{\rm a}/\tau_{\rm e}$ diagram.
The paper is organized as follows. We present an analytical model accounting for the dynamics of planets at resonances in Sec. 2. In Sec. 3 we perform N-body simulations and compare the results with the analytical predictions. In Sec. 4 we discuss several implications as well as potential caveats in our model. Finally, we summarize the key results in Sec. 5.
2 Dynamics of resonance trapping
In this section, we discuss the dynamical evolution of a two-planet pair in a gaseous protoplanetary disk after their convergent type-I migration. We consider two planets of masses $m_{\rm i}$ and $m_{\rm o}$ orbiting a central star of $M_{\star}$ on coplanar orbits. Their reduced masses and the mass ratio between the inner and outer ones are $\mu_{\rm i,o}{=}m_{\rm i,o}/M_{\rm\star}$ and $q{=}m_{\rm i}/m_{\rm o}$, respectively. For notations the subscripts ‘i’ and ‘o’ refer to the corresponding quantities of inner and outer planets.
We present the analytical criteria of resonance trapping and stability in Sec. 2.1 and Sec. 2.2, respectively. A unified picture is summarized in Sec. 2.3.
2.1 Resonance trapping criteria
The dissipative effect of planet-disk interaction can be parameterized by (Teyssandier & Terquem, 2014)
$\frac{1}{a}\frac{da}{dt}$ $=$ $-\frac{1}{\tau_{a}}-\frac{2e^{2}}{\tau_{e}},$ (1a)
$\frac{1}{e}\frac{de}{dt}$ $=$ $-\frac{1}{\tau_{e}},$ (1b)
where $\tau_{a}$ and $\tau_{e}$ are the orbital decay and eccentricity damping timescales, respectively. The orbital elements $a$, $n$, $e$, $\lambda$, $\varpi$, denote the semimajor axis, mean motion, eccentricity, mean longitude and longitude of pericenter of the planet.
Figure 1: Resonance crossing of a two-planet system with convergent migration. The plots provide the time evolution of the outer-to-inner planet period ratio (panel a), the two planets’ eccentricities (panel b), resonant angle (panel c) and phase curve (panel d). The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, the planet masses are $m_{\rm i}=1\ M_{\oplus}$, $m_{\rm o}=10\ M_{\oplus}$, and the migration parameters are $\tau_{a}=1.1\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=1.5\times 10^{3}$. The planets directly cross the $2$:$1$ resonance at $t{\simeq}6\times 10^{3}$ yr and their eccentricities simultaneously excite to $10^{-2}$ and rapidly decline afterward. The resonant angle of the inner planet jumps from the initial values to $\pi$ during the crossing and oscillates over $2\pi$ afterwards. In phase space, the inner planet is not captured by the non-zero resonant fixed point and ultimately evolves towards the origin.
For the $j$:$j-1$ order mean motion resonances (MMRs), the resonant angles and the time-averaged disturbing functions for the inner and outer planets are given by (Murray & Dermott, 1999) $\varphi_{\rm i}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i}$, $\varphi_{\rm o}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm o}$, and $R_{\rm i}{=}Gm_{\rm o}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+\alpha R_{\rm E})$, $R_{\rm o}{=}Gm_{\rm i}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+R_{\rm I}/\alpha^{2})$, where $\alpha{=}a_{\rm i}/a_{\rm o}$ is the two planets’ semi-major axis ratio, $R_{\rm D}^{\rm sec}$ and $R_{\rm D}^{\rm res}$ are the secular and resonant terms, $R_{\rm E}$ and $R_{\rm I}$ are the indirect part due to an external and internal perturber.
We assume small eccentricities of the planets, and expand the disturbing functions to the first order in $e_{\rm i}$ and $e_{\rm o}$, which gives $R_{\rm D}^{\rm sec}{=}0$, $R_{\rm D}^{\rm res}{=}e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$,$R_{\rm E}{=}-2e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, and $R_{\rm I}{=}-\frac{1}{2}e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, where $\delta_{\rm j,2}$ is the usual Kronecker symbol. Detailed expressions of $f_{\rm d,i}$ and $f_{\rm d,o}$ can be found in Table 8.1 of Murray & Dermott (1999).
Taking orbital decay and eccentricity damping into account, Lagrange’s equations can be written as
$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm i}}{2\tau_{a,\rm i}}+\frac{3n_{\rm i}e_{\rm i}^{2}}{\tau_{e,\rm i}},$ (2a)
$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm o}}{2\tau_{a,\rm o}}+\frac{3n_{\rm o}e_{\rm o}^{2}}{\tau_{e,\rm o}},$ (2b)
$\dot{e_{\rm i}}$ $=-\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\sin{\varphi_{\rm i}}-\frac{e_{\rm i}}{\tau_{e,\rm i}},$ (2c)
$\dot{e_{\rm o}}$ $=-\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}-\frac{e_{\rm o}}{\tau_{e,\rm o}},$ (2d)
$\dot{\varpi_{\rm i}}$ $=\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\frac{\cos{\varphi_{\rm i}}}{e_{\rm i}},$ (2e)
$\dot{\varpi_{\rm o}}$ $=\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\frac{\cos{\varphi_{\rm o}}}{e_{\rm o}},$ (2f)
where $f_{\rm d,o}^{{}^{\prime}}{=}f_{\rm d,o}-2\alpha\delta_{j,2}$. For instance, $f_{\rm d,i}{=}{-}1.19$, $f_{\rm d,o}^{{}^{\prime}}{=}0.43$ at the $2$:$1$ MMR.
When the planets are in MMRs, their orbital parameters reach the equilibrium state such that $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$, $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$. By substituting $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ into Laplace equation, Eqs. (2a) and (2b) can be combined as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (3)
$=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
where we assume that $\tau_{a}^{-1}{=}\tau_{a,\rm o}^{-1}-\tau_{a,\rm i}^{-1}$ is the relative orbital decay rate with respect to convergent migration, and $\tau_{e}{=}\tau_{e,\rm i}{=}\tau_{e,\rm o}$ represents the eccentricity damping time-scale. Substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$, the equilibrium eccentricities can be derived from Eq. (3) such that
$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4a)
$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4b)
where $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$.
Our results are analogous to Terquem & Papaloizou (2019), except for $\tau_{a}$ and $\tau_{e}$ as defined. In the limit of the mass of inner planet much lower than the outer one ($q{\ll}1$), the eccentricity of inner planet in Eq. (4a) can be simplified as
$e_{\rm i,eq}=\sqrt{\frac{1}{2j}\frac{\tau_{e}}{\tau_{a}}},$ (5)
in agreement with Eq. (24) in Huang & Ormel (2023). We note that due to differing signs of $f_{\rm d,i}$ and $f_{\rm d,o}^{{}^{\prime}}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ requires that $\sin{\varphi_{\rm i}}{>}0$ and $\sin{\varphi_{\rm o}}{<}0$ for $j$:$j-1$ MMRs. Considering $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$, we can derive $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}$, which implies $\varphi_{\rm i}-\varphi_{\rm o}{=}\pm\pi$. This means that the conjunctions between planet pair in a mean motion resonance always occur at either pericentre or apocentre.
Notably, the right-hand side of Eq. (3) represents the planet-disk interaction that leads convergent orbital migration between two planets, while the left-hand side of Eq. (3) represents their resonant gravitational interaction, keeping them away from getting closer. When the effect of planet-disk interaction is strong enough to overcome the resonant interaction, the planet pair cannot stay in resonance. Strong planet-disk interaction can be further classified into two regimes: fast migration and strong eccentricity damping. In the first regime the exchange rate of semimajor axis $\dot{a}/a$ in Eq. (3) is driven by the a-decay term ($\tau_{a}^{-1}{\gg}e^{2}\tau_{e}^{-1}$). In the second regime, however, the contribution of the e-damping term becomes comparable to that of the a-decay term. We will derive the resonance trapping conditions in the following subsections.
Figure 2: Resonance crossing of a two-planet system with convergent migration. The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=10^{5}yr$ and $\tau_{a}/\tau_{e}=6\times 10^{2}$. The planets are trapped in the $2:1$ resonance at $t{\simeq}10^{4}$ yr and their eccentricities simultaneously excite to $2\times 10^{-2}$, which remain stable thereafter. The resonant angle of the inner planet is excited when the period ratio evolves to 2:1, subsequently oscillating and gradually converging to the resonant equilibrium value, which it maintains. In phase space, the inner planet is captured by the non-zero resonant fixed point and subsequently evolves in conjunction with this fixed point, reflecting the static equilibrium state of the resonance..
2.1.1 Slow migration
When the planets undergo fast convergent migration, they directly bypass the resonance before the eccentricities reach to the equilibrium values. In this case $\dot{a}/a$ is driven by $\tau_{a}$ but not $\tau_{e}$, and Eq. (3) can be rewritten as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}},$ (6)
where the e-damping term is ignored. Here the eccentricities are purely excited by the resonant term, which can be deduced from the classical pendulum model (see appendix Sec. Appendix A: The derivation of the excited eccentricity in the classical pendulum model for derivation)
$e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}=-\frac{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}.$ (7)
Combing Eqs. (6) and (7), the resonance trapping criterion of slow migration can be obtained as
$\tau_{a}>\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}.$ (8)
We further simplify the above criterion as
$\begin{cases}{\displaystyle\tau_{a}>\frac{1}{2n_{\rm i}(3(j-1))^{1/3}}\frac{1}{(\mu_{\rm o}\alpha f_{\rm d,i})^{4/3}}\ \sim\mu_{\rm o}^{-4/3}}\ \ \hfill\mbox{ at $q\ll 1$},\\ {\displaystyle\tau_{a}>\frac{1}{2n_{\rm o}(3j)^{1/3}}\frac{1}{(\mu_{\rm i}f_{\rm d,o}^{{}^{\prime}})^{4/3}}\ \sim\mu_{\rm i}^{-4/3}}\ \ \ \hfill\mbox{at $q\gg 1$}.\\ \end{cases}$ (9)
This indicates that the threshold migration rate increases with the planet mass. Intuitively, the planet resonant interaction becomes stronger as their masses increase. Therefore, the planets require faster migration in order to break the resonance.
2.1.2 Weak eccentricity damping
When the planets migrate relatively slow, they have time to reach the equilibrium state. Substituting the equilibrium eccentricity into the right side of Eq. (3), it turns out that the magnitude of e-damping term is comparable to that of the a-decay term. In this circumstance, the e-damping term becomes a key factor for resonance trapping.
Considering the limit of $\sin{\varphi_{\rm i}}{=}1$ and $\sin{\varphi_{\rm o}}{=}-1$, Eq. (3) can be rewritten as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$ (10)
where the left-hand side of the planetary interaction maintains a finite magnitude at resonances (${\propto}\mu ne$). If eccentricity damping is very strong, the orbital circularization leads large $\dot{a}/a$,, and the resonance can no longer hold. One example of resonance crossing is shown in Figure 1. Hence, resonance trapping requires weak e-damping. Substituting Eq. (10) with equilibrium values of $e_{\rm i}$ and $e_{\rm o}$ from Eqs. (4a) and (4b), we obtain the resonance trapping criterion in the weak e-damping regime as
$\tau_{e}\tau_{a}>\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}.$ (11)
In the limit of $q{\ll}1$, the trapping criterion can be simplified as
$\tau_{e}\tau_{a}>\frac{1}{2j(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i})^{2}},$ (12)
which is in consistent with Eq. (25) of Huang & Ormel (2023).
Altogether, the resonance trapping requires both relatively weak eccentricity damping and slow migration. Taking Eqs. (11) and (8) together, we formulate the necessary condition for $\tau_{a}$ and $\tau_{e}$ as
$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}].\end{cases}$ (13)
Figure 3: Resonance crossing of a two-planet system with convergent migration. The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=4\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=5\times 10^{2}$. The planets are trapped in the $2:1$ resonance at $t{\simeq}3\times 10^{4}$ yr and their eccentricities first excite to $2\times 10^{-2}$, and starts to exhibit significant oscillations around its equilibrium value. The resonant angle of the inner planet remains at the equilibrium value and begins to oscillate after some time. In phase space, the inner planet initially evolves to a non-zero eccentricity resonant fixed point and subsequently starts oscillating around it. Over time, the system settles into a limit cycle around the fixed point, reflecting the dynamic equilibrium state of the resonance.
2.2 Post-evolution after resonant trapping
Here we assume that the planet pairs have satisfied the resonance trapping conditions in Sec. LABEL:sec:riteria and study their following dynamical evolution. The goal is to explore under which configuration the resonance state is temporary or long-term stable.
Our approach follows the analysis outlined in Section 2 of Xu & Lai (2017) (also see Chapter 8.8 of Murray & Dermott 1999). We first focus on the situation that the mass of the inner planet is much lower than the outer one ($m_{\rm i}{\ll}m_{\rm o}$, $q{\ll}1$). In the absence of planet-disk interactions, the system’s Hamiltonian $H$ near the $j$:$j-1$ MMR can be expressed as
$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (14)
where
$\varphi$ $=$ $j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i},$ (15a)
$\Phi$ $=$ $[\frac{3(j-1)^{2}}{f_{\rm d,i}\mu_{\rm o}\alpha_{\rm 0}}]^{2/3}\frac{1-\sqrt{1-e_{\rm i}^{2}}}{2}\sim\mu_{\rm o}^{2/3}e_{\rm i}^{2},$ (15b)
$\eta$ $=$ $[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}]^{2/3}[(j-1)\alpha_{\rm 0}^{-2/3}-j\alpha_{\rm 0}^{5/6}].$ (15c)
The first two parameters are a pair of conjugate coordinate and momentum, where $\varphi$ is the resonant angle and $\Phi$ represents the angular momentum deficit (AMD). The third parameter, $\eta$ is the dimensionless that quantifies the deviation level from the exact resonance. The orbital period ratio of the two planets is the exact integer ratio at $\eta{=}0$. The initial semi-major axis ratio of the two planet $\alpha_{\rm 0}\equiv\alpha[1+(j-1)e_{\rm i}^{2}]$ is the conjugate momentum of a fast angle, which can be derived from Eq.8-92 of Murray & Dermott (1999). We note that in the absence of planet-disk interactions, the parameters $\eta$ and $\alpha_{0}$ remain constant.
As we will demonstrate later, the effects of planet-disk interaction ($\tau_{\rm a},\tau_{\rm e}$) determines the stability of the planet pair in resonance.
2.2.1 stability/overstability criterion
When accounting for the planet-disk interaction, the equations of motion of the system can be obtained from
$\dot{\Phi}{=}-\frac{\partial H}{\partial\varphi}+\frac{\partial\Phi}{\partial e}\frac{de}{dt},\ \ \ \dot{\varphi}{=}\frac{\partial H}{\partial\Phi}.$ (16)
By incorporating the orbital decay and eccentricity damping terms in Eq.(1), Eq. (16) can be derived as
$\dot{\Phi}$ $=$ $\Phi^{1/2}\sin{\varphi}-\frac{2}{\tau_{e}}\Phi,$ (17a)
$\dot{\varphi}$ $=$ $-\eta-2\Phi+\frac{1}{2\sqrt{\Phi}}\cos{\varphi},$ (17b)
$\dot{\eta}$ $=$ $\frac{\partial\eta}{\partial\alpha_{0}}\frac{d\alpha_{0}}{dt}{=}-\frac{\beta}{\tau_{a}}+\frac{2(p+2j-2)\Phi}{(j-1)\tau_{e}},$ (17c)
where
$\beta\equiv|\frac{\partial\eta}{\partial\ln\alpha_{\rm 0}}|=[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}]^{2/3}\frac{3}{2}(j-1)^{5/9}j^{4/9},$ (18)
and $\Phi{=}\beta[(j-1)e_{\rm i}^{2}/2]$.
The system reaches an equilibrium state at resonances when $\dot{\Phi}{=}0$, $\dot{\varphi}{=}0$, $\dot{\eta}{=}0$. Thus, the resonant parameters, to the lowest order, can be written as
$\Phi_{\rm eq}$ $=$ $\frac{(j-1)\beta\tau_{e}}{4j\tau_{a}},$ (19a)
$\varphi_{\rm eq}$ $\simeq$ $\frac{2}{\tau_{e}}\Phi_{\rm eq}^{1/2},$ (19b)
$\eta_{\rm eq}$ $\simeq$ $\frac{1}{2\sqrt{\Phi_{\rm eq}}}-2\Phi_{\rm eq}.$ (19c)
We then use the linear perturbation method to calculate these quantities ($\Phi{=}\Phi_{\rm eq}+\delta\Phi$ etc.). The corresponding linearized equations read
${\begin{bmatrix}\delta\dot{\Phi}\\ \delta\dot{\varphi}\\ \delta\dot{\eta}\end{bmatrix}=\begin{bmatrix}-\tau_{e}^{-1}&\Phi_{\rm eq}^{1/2}&0\\ -(0.25\Phi^{-3/2}+2)&-\tau_{e}^{-1}&-1\\ 4j\tau_{\rm e}^{-1}(j-1)^{-1}&0&0\end{bmatrix}\begin{bmatrix}\delta{\Phi}\\ \delta{\varphi}\\ \delta{\eta}\end{bmatrix}}.$ (20)
We assume that the solutions have the following forms $\delta\Phi,\delta\varphi,\delta\eta\propto e^{\lambda t}$ and $\lambda$ is a complex number. The characteristic equation can be written as
$\lambda^{3}+\frac{2}{\tau_{e}}\lambda^{2}+\left(2\Phi_{\rm eq}^{1/2}+\frac{1}{4\Phi_{\rm eq}}\right)\lambda+\frac{\beta}{\tau_{a}\Phi_{\rm eq}^{1/2}}=0.$ (21)
We provide a general approach to solve the real part solutions of Eq. (21) as folows. Suppose that a cubic equation has a form of $A\lambda^{3}+B\lambda^{2}+C\lambda+D{=}0$ where $A,B,C,D{>}0$. Assuming that $\lambda{=}\lambda_{\rm r}\pm i\lambda_{\rm i}$, it can be proved that when $D{<}BC/A$, the real part of eigenvalues $\lambda_{r}{<}0$. This consequently means that the trapping is stable where the trajectory of Hamiltonian $H$ converges to the fixed point (see an illustration in Figure 2). Substituting $A,B,C,D$ with Eq. (21), we obtain our stability criterion as
$\frac{\tau_{a}}{\tau_{e}}>\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{2/9}j^{4/9}}{2j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (22)
Our result is consistent with Eq. (26) of Goldreich & Schlichting (2014), where they constructed different forms of conserved quantities for derivations (see their Eq. 9).
Our analysis can also be extended to the case when the mass of the outer planet is much lower than that of the inner one ($m_{\rm i}{\gg}m_{\rm o}$, $q{\gg}1$). In such a circumstance, we find that the real part of eigenvalues of characteristic equation $\lambda_{\rm r}{<}0$ always holds. Importantly, this means that once the planet pair gets trapped in the resonance, the trapping is always stable as long as $m_{\rm i}{\gg}m_{\rm o}$.
Figure 4: Level curves of the Hamiltonian for different $\eta$, plotted in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. The black (red) dots mark the stable (unstable) fixed points, the thick black line in panel (b) marks the separatrix. Each curve represents the conserved Hamiltonian in the absence of dissipitive disk forces. The color of the lines represents the values of the Hamiltonian. When considering the planet-disk interaction, during the convergent migration of the planet, H increases, which corresponds to a color shift from purple and green to yellow. When overstability occurs, the perturbations grow, causing the curve to expand, and H decreases, corresponding to a color shift from yellow and green to purple.
2.2.2 Escape criterion
When planets are in resonances and under the stable criterion, the resonant quantities ($\phi$,$\Phi$,$\eta$) remain fixed with a small amplitude liberation. Here we attempt to answer what occurs if the above stability criterion is not satisfied. We define such a state as overstablity. In this case, the perturbations grow exponentially over time (no longer in the linear regime). Whether the overstability leads planets directly escape or still stay in resonance can be further analyzed by the phase space topology.
Figure 4 shows the Hamiltonian curves for different $\eta$ values in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. We note that $\eta$ represents the deviation from the exact resonance and the distance to the origin corresponds to the excited eccentricity strength ($\sqrt{X^{2}+Y^{2}}{=}\sqrt{\Phi}{\sim}e$).
Importantly, the value of $\eta$ sets the number of fixed points, and therefore determines the subsequent dynamical evolution. The fixed points can be obtained from $\partial H/\partial\varphi{=}\partial H/\partial\Phi{=}0$ (also see Eq.8-135 of Murray & Dermott 1999). It can be shown that, for $\eta{>}-3/2$, there is only one stable fixed point (black dot in Figure 4), whereas for $\eta{<}-3/2$, there are two stable fixed points and one unstable point (red dot in Figure 4).
In the overstable circumstances, the perturbation near the equilibrium grows significantly. When the equilibrium value $\eta_{\rm eq}{>}-3/2$, the system’s Hamiltonian curve expands as eccentricities excited during resonant capture. However, the trajectory still bounds to the only fixed point. As a result, the system’s eccentricity, period ratio, and resonant angle librate near the equilibrium. Following Goldreich & Schlichting (2014) and Xu & Lai (2017), we refer to this state as an overstable trap, which is also equivalent to circumstance described as the limit cycle in Afkanpour et al. (2024). One example of overstable trap can be found in Figure 3.
The separatrix occurs when $\eta_{\rm eq}{<}-3/2$ and the topology completely changes, as shown in Figure 4. In this case, the system’s Hamiltonian curve originally sits near the right stable fixed point. As the curve expands, it cross the separatrix and enter the circulation zone centered around the left stable fixed point (from the right to the left black dot in Figure 4). After then, the curve continuously shrinks towards the left black fixed point. As such, the eccentricity decreases, and the weaker resonant interaction cannot keep the planet pair in resonance anymore. The planets’ period ratio departs from the exact commensurability, and the system eventually escapes from the resonance. One example of escape is shown in Figure 5.
Based on Eq. (19c), $\eta_{\rm eq}{<}-3/2$ requires $\Phi_{\rm eq}{=}(j-1)\beta\tau_{\rm e}/(4j\tau_{a}){>}1$, which is corresponding to
$\frac{\tau_{a}}{\tau_{e}}<\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{8/9}j^{4/9}}{8j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (23)
The above escape criterion agrees with Eq. (28) of Goldreich & Schlichting (2014).
It is worth noting that this escape condition requires the system is already in the overstable regime. In the case where the mass of the outer planet is much lower than that of the inner one ($q\gg 1$), the resonance is always stable. As such, the escape does not occur even when $\eta_{\rm eq}{<}-3/2$.
Taking together, resonance trapping is always stable when the mass of the inner planet is much higher than that of the outer one ($q{\gg}1$). However, when the mass of the outer planet is much higher than that of the inner one ($q{\ll}1$), the outcome can be classified into three configurations: stable trap, overstable trap, or escape. We further extend the analysis to arbitrary mass ratios $q$ and derive the corresponding criteria as follows:
$\begin{cases}{\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}].\end{cases}$ (24)
where
$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (25)
$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (26)
Mention that $h(q)=\tau_{a}/\tau_{e}(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2})$, $f(q)=8(j-1)[1-(e_{\rm o,eq}/e_{\rm i,eq})^{2}]$. Eq. (24) reduces to Eqs. (22) and (23) in the limit $q\ll 1$. We show the detailed derivation in the Appendix B.
Figure 5: Resonance crossing of a two-planet system with convergent migration. The initial starting from the outer-to-inner planet period ratio is $2.2$ (slightly beyond the $2$:$1$ MMR), and migration parameters are $\tau_{a}=2.5\times 10^{5}$ and $\tau_{a}/\tau_{e}=10^{2}$. The period ratio between the inner and outer planets exhibits small oscillations upon entering resonance and ultimately decreases further away from the 2:1 ratio. The eccentricity of the inner planet undergoes significant oscillations around the resonant equilibrium value after resonance is established, eventually damping to a very small value. The resonant angle of the inner planet maintains large oscillations upon entering resonance and eventually diverges. In phase space, the inner planet initially oscillates around the resonant fixed point, but the oscillation range increases until it is ultimately captured by the origin.
Figure 6: The theoretical distribution of the four kinds of evolution near 2:1 MMR depending on different $\tau_{a}$ and $\tau_{a}/\tau_{e}$. When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blue region, the two planets pass through the resonance because the eccentricity damping is too strong. When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blown region, they can’t be trapped into the resonance since they migrate inward too fast. The green region indicates that the planets can maintain a stable resonance. The yellow area represents that although the resonance trapping is overstable, the planets can still maintain a 2:1 period ratio. The purple region indicates that the planets will eventually leave the 2:1 MMR position and migrate inward after overstability. The left panel shows the distribution when the outer planet is more massive than the inner one, and the right panel displays the distribution when the inner planet is more massive than the outer one.
2.3 A unified picture
We summarize different outcomes considering the dynamics evolution of planet pairs that undergo convergent migration. The illustrated runs from the numerical simulations in each regime are also presented.
  1. No trap (Figure 1): The system cannot reach the resonant equilibrium state and the planet pair directly bypasses the resonance. The planets’ eccentricities increase and their period ratio temporarily holds during the resonant passage. However, after rapid crossing the resonance, the eccentricity drops and the resonant angle varies from 00 to $2\pi$.
  2. Stable trap (Figure 2): The planets are locked in resonance with statically stable orbits. The planets’ period ratio, eccentricities, and resonant angle remain at the equilibrium values after capture. Meanwhile, the trajectory of the system converges towards a fix point in the phase space.
  3. Overstable trap (Figure 3): The planets are locked in resonance with dynamically stable orbits. Their eccentricities and resonance angle initially converge to the equilibrium values. After a while, the eccentricities and resonance angle liberate with increasing amplitudes. This state is nevertheless stable and the period ratio of planets remains constant after capture. In the phase space, the system’s trajectory initially approaches the fixed point, then begins to oscillate, and ultimately settles into a limit cycle around that point.
  4. Escape (Figure 5): The planets are locked in resonance with only temporarily stable orbits. Their period ratio, eccentricities and resonance angle initially converge to the equilibrium values. The planets’ eccentricities and resonance angle quickly oscillate with increasing amplitudes. This state becomes unstable and the planets eventually leave resonance. As such, the planets’ eccentricities decrease, and the period ratio continues to drop, and the resonant angle changes from liberation to circulation. In the phase space, the system’s trajectory initially approaches the fixed point; however, the oscillation gradually grows until it intersects the separatrix. The trajectory ultimately settles towards the coordinate origin.
We point out that although several literature studies have explored this issue in depth (Huang & Ormel, 2023; Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Batygin, 2015), none have formulated a comprehensive model. Here we demonstrate a unified picture in Figure 6. The outcome is given in the $\tau_{a}$-$\tau_{e}$ parameter space. We describe how $\tau_{\rm a}$ and $\tau_{\rm e}$ influence the outcomes as follows.
If convergent migration is too rapid, the planets directly bypass the resonance without reaching the equilibrium state. As migration slows down, the effect of eccentricity damping becomes significant. Strong e-damping has two consequences: the planet resonant interaction gets weaken and semimajor axis decays due to orbital circularization. When the orbital decay surpasses the resonant term, the planets will no longer remain in resonance. In this regard, planets cannot get capture when $\tau_{\rm e}$ and/or $\tau_{\rm a}$ are short. No trap occurs in the upper and left part of Figure 6.
When planets are trapped in resonance, the stability is related to the degree of eccentricity excitation. The damping effect ($\tau_{\rm e}$) reduces the planet’s eccentricity, leading the system stable. On the other hand, convergent migration ($\tau_{\rm a}$) brings planets closer together, in turn exciting their eccentricities. Consequently, the relative eccentricity damping $\tau_{\rm a}/\tau_{\rm e}$ is the key parameter affecting the stability. As can be seen from top to bottom of Figure 6, system tends to be more unstable as $\tau_{\rm a}/\tau_{\rm e}$ decreases.
For the small inner planet case shown in Figure 6, if the relative damping $\tau_{\rm a}/\tau_{\rm e}$ is sufficiently strong, the resonant trapping is statically stable, and the final state of the system stays at a fixed point in the phase space (stable trap regime). As the relative damping weakens (Eq. 22), the fixed point becomes unstable, and the system’s trajectory begins to oscillate around the fixed point. It ultimately settles into a stable limit cycle (overstable trap). The limit cycle expands as $\tau_{\rm a}/\tau_{\rm e}$ continues to decrease. When the trajectory crosses the separatrix (Eq. 23), the system becomes dynamically unstable. In this regime, the oscillations grow over time and planets eventually leave resonance (escape). In contrast, for the case of a small outer planet shown in Figure 6, all traps are stable.
Figure 7: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $0.1$. Blue points indicate cases where the planets are not trapped by the resonance, green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate that the planets eventually escape from the resonance. The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{a}\sim 10^{5}\ \mathrm{yr}$ and $\tau_{a}/\tau_{e}\sim 3\times 10^{2}$. While the analytical model predicts overstable trapping in this regime, the simulations show escape from resonance. We hypothesize that as migration becomes faster, the perturbations grow stronger and become nonlinear, leading to amplified oscillation amplitudes and an increased likelihood of escaping from resonance.
Figure 8: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $10$. The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$. While the analytical model predicts resonance trapping in this regime, the simulations show resonance crossing. We hypothesize that the discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect.
Figure 9: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q=1$. We compare our criteria with previous studies.We cite the trapping criterion from Huang & Ormel (2023) and the stable and escape criteria from Goldreich & Schlichting (2014). The solid lines represent our criteria, the dash lines represent the criteria cited from previous studies. In the case of equal-mass planets, it is evident that our criteria align more closely with the simulation results.
3 Simulations
In this section, we conduct numerical simulations for $2:1$ resonance to verify the different regimes and criteria analytically derived in the previous section. The numerical setup is described in Sec. 3.1. The comparison between simulations and analysis at two extreme mass ratios is presented in Sec. 3.2, and the extrapolation for any mass ratio is discussed in Sec. 3.3.
3.1 Numerical setup
We use the open-source N-body code REBOUND with the WHfast integrator (Rein & Liu, 2012). The simulations are confined in conplanar circumstances, and the accelerations for orbital decay and eccentricity damping are given by
${a_{\rm m}}=-\frac{\textbf{v}}{2\tau_{a}},\ \ {a_{\rm e}}=-2\frac{(\textbf{v}\cdot\textbf{r})\cdot\textbf{r}}{r^{2}\tau_{e}},$ (27)
which are implemented through REBOUNDx (Tamayo et al., 2020).
Initially, we place the inner and outer planets at $1$ and $1.7$ au, respectively, making them slightly away from the $2$:$1$ MMR. Both planets have circular orbits with random phase angles. To mimic convergent migration, we only let the outer planet undergo inward migration on a time-scale of $\tau_{\rm a}$, whereas the eccentricities of both planets are damped on a time-scale of $\tau_{\rm e}$. Simulations are terminated at $t{=}2\tau_{\rm a}$, and the final outcomes of the planets in each run are determined as described in Sec. 2.3.
We consider three cases for the masses of the planet-pair: small inner planet ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$), small outer planet ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), equal-mass planets ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$). We vary different sets of $\tau_{\rm a}$ and $\tau_{\rm e}$ in each simulations.
3.2 Comparison at two extreme mass ratios
The dynamical outcomes of planet pairs that undergo convergent migration are shown in Figure 6. The results are dependent on the mass ratio of the planets and disk properties ($\tau_{\rm a}$ and $\tau_{\rm e}$). Basically four regimes are presented when the inner planet is lower than that of the outer one. The illustrated runs in each regime are provided in Figures 1,2,3 and 5.
Figure 8 shows the numerical results obtained from an illustrated case where the mass of inner planet is ten times lower than that of the outer one ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$). The numerical results shown in Figure 8 can directly be compared with the schematic presented in Figure 6.
We find that the simulations agree pretty well with the analytical estimates. No planets lead resonant trap (blue dots) in the upper left part of Figure 8, when the weak damping (upon the green line) and slow migration (left of the brown line) criteria break. When those trapping criteria are satisfied and as $\tau_{\rm a}/\tau_{\rm e}$ decreases, the simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes. These results are indeed separated by the two analytical criteria proposed in Sect. 2 (yellow and purple lines).
However, there is one point in Figure 8 that simulations do not fully match the analysis, occurring at $\tau_{a}{\sim}10^{5}\ {\rm yr}$ and $\tau_{a}/\tau_{e}{\sim}3\times 10^{2}$. The theory predicts the planets in overstable trap regime whereas in simulations planets actually leave the resonance. We speculate that the reason is that as migration becomes faster, the perturbations intensify and become nonlinear, resulting in further increasing oscillation amplitudes around the equilibrium value. Consequently, the system becomes more likely to cross the separatrix and escape resonance.
Figure 8 demonstrates simulation results from the case when the mass of inner planet is ten times higher than that of the outer one ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), serving as a comparison to the schematic illustrated in Figure 6. Again, no trap (blue dots) occurs in the upper and left parts of Figure 8, when the weak damping and slow migration criteria are not satisfied. When these criteria are fulfilled, all traps correspond to stable trap (green dots).
We note that the no-trap regime occupies a large parameter space than theoretically expected. This is evident at $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$, where the trapping criteria are fulfilled but the planets are not in resonance. This discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect (Eq. 6). When the eccentricity damping is considered, the interaction between the planets is further suppressed (lower $e$ in left-hand side of Eq. 3), leading to a longer migration timescale (right-hand side of Eq. 3). Consequently, in reality planets can bypass the resonance with slower migration rates compared to the simplified estimation from Eq. 6.
In conclusion, our analytical predictions align well with simulations in the circumstances of planet pairs with extreme mass ratios.
3.3 Comparison at arbitrary mass ratios
Notably, Huang & Ormel (2023) derived the weak damping trapping criterion at the small mass ratio limit of $q{\ll}1$. Goldreich & Schlichting (2014) investigated when the trap becomes overstability and/or escape in the same limit. In this work we derived the general criteria for trap and stability that are well suited to the arbitrary mass ratio (Eqs. 13 and 24). In this subsection we show the simulation comparison, whereas the detailed derivation of the $q$-dependent overstabity and escape criteria is provided in Appendix B.
Figure 9 illustrates the simulation results for the equal-mass planet case ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$) and compares our criteria with those from Huang & Ormel (2023) and Goldreich & Schlichting (2014). The solid lines represent our analytical criteria, dashed lines correspond to criteria from previous studies, and colored dots depict the simulation results. We find that the new criteria clearly depict a better agreement with numerical results.
Figure 10: The plots illustrate the dependence of resonance trapping criteria on the semi-major axis (panel a) and the stellar mass (panel b) for a two-planet system, based on the minimum solar nebular model (MMSN). The planet masses are $m_{\rm i}=1\ M_{\oplus}$ and $m_{\rm o}=10\ M_{\oplus}$. In panel a, the stellar mass is fixed at $M_{\star}=1\ M_{\odot}$, while in panel b, the semi-major axis of the inner planet is fixed at $a_{\rm i}=1\ au$. The brown line represents the slow migration trapping criterion (Eq. 8), below which migration is too fast for resonance capture. Regions above the brown line indicate where capture is possible. The color gradient represents the strongest eccentricity damping $\tau_{a0}/\tau_{e0}$ that allows resonant capture (Eq. 11), with lighter colors meaning the planets can be in resonances with stronger eccentricity damping. In panel (a), as the semi-major axis increases, the critical value of $\tau_{a0}$ remains constant, while $\tau_{a0}/\tau_{e0}$ slightly increases. This suggests that planets are more easily trapped in resonances under stronger eccentricity damping as the semi-major axis increases. In panel (b), as the stellar mass decreases, the critical value of $\tau_{a0}$ decreases, and $\tau_{a0}/\tau_{e0}$ increases. This indicates that planets are more easily trapped in resonances under stronger eccentricity damping and faster migration as the stellar mass decreases.
4 Discussion
In this section, we connect our criteria to disk parameters in Sec. 4.1 and analyze how the criteria depend on the planet location and the mass of stellar host. Additionally, we investigate resonance trapping in the context of divergent migration in Sec. 4.2 and discuss several caveats of our work in Sec. 4.3.
4.1 Dependence on semi-major axis and stellar mass
Here we explore how the resonant trapping conditions depend on the radial distance and stellar mass. The detailed type I migration formulas for low-mass planets are given by $\tau_{a}^{-1}{\sim}\mu\mu_{\rm d}h^{-2}n$ and $\tau_{e}^{-1}\sim\mu\mu_{\rm d}h^{-4}n$, where $\mu_{\rm d}{=}\Sigma_{\rm d}a^{2}/M_{\rm*}$ is the mass ratio between disk and star, and $\Sigma_{\rm d}$ and $h$ are the disk surface density and aspect ratio, respectively.
Following the minimum mass solar nebula (MMSN, Hayashi 1981) model, the disk surface density and aspect ratio can be written as
$\Sigma_{\rm d}=$ $1700\left(\frac{a}{\rm 1\ au}\right)^{-3/2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)\rm\ g/cm^{2},$ (28)
$h=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{L_{\star}}{1\ L_{\odot}}\right)^{1/8}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/2}$
$=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/4},$ (29)
the latter of which $L_{\star}{\propto}M_{\star}^{2}$ is assumed for the pre-main sequence stars. By substituting $\Sigma_{\rm d}$ and $h$ into $\tau_{a}$ and $\tau_{e}$, we obtain their dependence on the semi-major axis and stellar mass as
$\tau_{a}={\tau_{a0}}\left(\frac{a}{\rm 1\ au}\right)^{3/2},\ \ \tau_{e}={\tau_{e0}}\left(\frac{a}{\rm 1\ au}\right)^{2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{1/2},$ (30)
where $\tau_{a0}$ and $\tau_{e0}$ represent the values of $\tau_{a}$ and $\tau_{e}$ at $1$ au around a solar-mass host. By inserting $\tau_{a0}$ and $\tau_{e0}$ into Eq. (13), we can separate the dependence of resonance trapping on semi-major axis and stellar mass.
Figure 10 shows the a-dependency for resonance trapping, where $M_{\star}{=}1\ M_{\odot}$ and $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$. The region below the brown line (slow migration criterion) indicates that migration is too fast for resonance capture, while the region above suggests that capture is possible, which depends on $\tau_{a0}/\tau_{e0}$. The color in Figure 10 refers to the critical $\left(\tau_{a0}/\tau_{e0}\right)$ that allows resonant capture. A lighter color means the planets can be in resonances with higher $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ (stronger e-damping). As the semi-major axis increases, $(\tau_{a0})_{\rm slow,mig}$ remains constant and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ slightly increases. This means that planets would be more easily trapped at resonances under stronger e-damping as $a$ increases. Nevertheless, the capture reveals an overall weak semi-major axis dependence.
Figure 10 shows $M_{\star}$-dependency for resonance trapping. We fix the inner planet at $1\ au$ and vary with the stellar mass. From the figure, as the stellar mass decreases, $(\tau_{a0})_{\rm slow,mig}$ decreases, and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ increases. This implies that planets can be trapped under a wider range of $\tau_{a}$ and $\tau_{e}$ parameter space. In other words, planets are more likely to be capture into resonance as the stellar mass decreases.
Based on the MMSN model, we find that planets are more likely to get capture into resonance with increasing semi-major axis and/or decreasing stellar mass. However, we note that these results are considerably sensitive to the adopted disk model. The interpretation should be taken with caution.
4.2 Divergent migration
We have shown in the previous sections that the convergent planet migration coupled with eccentricity damping can result in resonance trapping. Literature theoretical work suggested that planets cannot be captured in resonance when the direction of their migration is divergent (Wong & Lee, 2024). Based on our analysis, here we also provide a simple argument that why resonance trapping cannot occur through divergent migration.
When the planet pair undergoes the divergent migration, the equilibrium state $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ (if it exists) can be written as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (31)
$=-\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
where the relative orbital decay term ($-\tau_{a}^{-1}$) is negative compared to Eq. (3). In order to derive equilibrium eccentricity, we repeat the procedure outlined in Sec. 2.1 by substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$. Then Eq. (31) reads
$\frac{e_{\rm i}^{2}}{\tau_{\rm e}}(1+q\sqrt{\alpha})[j+(j-1)\frac{f_{\rm d,o}^{{}^{\prime}2}}{f_{\rm d,i}^{2}}q\sqrt{\alpha}]=-\frac{1}{2\tau_{a}}.$ (32)
Noticeably, the left-hand side of this equation is positive, whereas the right-hand side is negative. This indicates that there is no equilibrium solution in the divergent migration circumstance. We also conducted numerical simulations and consistently found that even a small amplitude of divergent migration is sufficient for planets to cross the resonance.
4.3 Caveats
Here discuss a few limitations in our model. First, for simplicity we treat $\tau_{\rm a}$ and $\tau_{\rm e}$ as free but constant parameters and two planets have the same eccentricity damping timescales. However, in reality $\tau_{\rm a}$ and $\tau_{\rm e}$ depend on the gas disk mass, planet mass and orbital distance. For instance, $\tau_{\rm a}$ and $\tau_{\rm e}$ decrease as planet mass increases, which scale with $\mu^{-1}$. The two trapping criteria exhibit steeper decreasing functions of $\mu$ such that $\tau_{\rm a,\ \text{slow\ mig}}{\propto}\mu^{-4/3}$, and $(\tau_{\rm a}\tau_{\rm e})_{\rm\text{weak\ damp}}{\propto}\mu^{-2}$. This suggests that planets become more easily trapped as their mass increases. On the other hand, since the planets can grow their masses at different rates, the planet mass ratio could be time varied. For instance, if the inner planet is initially more massive than the outer planet ($q{\gg}1$), the resonance trapping is the stable regime. When the outer planet grows and the mass ratio decreases, the system may transition into an overstable trap or even escape. Moreover, both hydrodynamic simulations and analytical work have shown that the migration and eccentricity damping rates differ from those in the near circular case when planets attain high eccentricities (Cresswell & Nelson, 2008; Ida et al., 2020). Future studies are required to investigate how these factors influence our analysis.
Second, we only consider a static disk. The configuration may change when disk dissipation is taken into account. In this case, the migration slows down as the gas density drops, making the resonance trapping easier. However, for stability/overstability and escape criteria, the situation is less straightforward. This is because $\tau_{\rm a}/\tau_{\rm e}$ depends on the aspect ratio of the disk but not the surface density. The disk thermal structure relies on the heating mechanism and therefore a specific disk model. As such, our framework cannot predict whether the planet pairs in the stable or overstable regimes will remain trapped as gas disk dissipates.
In addition, the disk is assumed to be laminar here. In a more realistic turbulent disk, density fluctuations generated by turbulence can result in additional stochastic torques (Goldberg & Batygin, 2023; Rein, 2012). This effect cannot be captured by our simplified approach that assumes static migration and eccentricity damping timescales. However, the outcomes of resonance trapping may be altered by considering the stochastic torques, which is a key focus of our future work.
5 Conclusion
We have built an analytical framework suited to study the resonant trapping, stability, and escape conditions for a two-planet pair that undergoes the convergent type I migration in a gaseous disk. The $2$:$1$ first order mean motion resonance is specifically investigated and the key outcomes are illustrated in Fig 6.
Planets can be trapped in resonances under the circumstances of slow migration and weak damping (Eqs. (8) and (11)). When the mass of inner planet is comparable or less massive than the outer planet ($q{\lesssim}1$), after capture the stability weakens as $\tau_{\rm a}/\tau_{\rm e}$ decreases. The system transitions from a stable trap to an overstable trap (Eq. 22), and finally to escape (Eq. 23). This corresponds to the resonant states changing from statically stable to dynamically stable, and ultimately to dynamically unstable. In contrast, when the mass of inner planet is much more massive than the outer planet ($q{\gg}1$), the traps remain permanently stable.
We provide the general analytical criteria for planets with an arbitrary mass ratio $q$ in Sec. 2. The corresponding formulas are also summarized as follows.
$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (33)
where
$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (34)
$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (35)
We validate the theoretical predictions with numerical N-body simulations (Figures 8, 8 and 9). Furthermore, within the framework, we show that resonance trapping cannot occur when planets migrate divergently. We also explored the dependence of the semi-major axis and the stellar mass based on minimum mass solar nebular (MMSN) model, which indicates planets are more likely to be resonance-trapped with increasing semi-major axis and decreasing stellar mass.
Acknowledgements.
This work is supported by REBOUND (Rein & Liu, 2012) and REBOUNDX (Tamayo et al., 2020)
Appendix A: The derivation of the excited eccentricity in the classical pendulum model
When the planets are in resonance, the second time derivative of the resonant angle $\varphi_{\rm i}$ and $\varphi_{\rm o}$ can be written as
$\ddot{\varphi_{\rm i}}{=}\ddot{\varphi_{\rm o}}{=}j\dot{n_{\rm o}}+(1-j)\dot{n_{\rm i}},$ (A1)
where the contribution from $\dot{\varpi_{\rm i}}$ and $\dot{\varpi_{\rm o}}$ are neglected. With no additional effects from the gaseous disk such as orbital decay and eccentricity damping, Lagrange’s equations can be written as
$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}),$ (A2a)
$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}),$ (A2b)
$\dot{e_{\rm i}}$ $=-\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\sin{\varphi_{\rm i}},$ (A2c)
$\dot{e_{\rm o}}$ $=-\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}.$ (A2d)
Substituting $\dot{n_{\rm o}}$ and $\dot{n_{\rm i}}$ in Eq. (A1) with Eqs. (A2a) and (A2b), and combing Eqs. (A2c) and (A2d), we obtain
$\ddot{\varphi}{=}3[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A3)
$\dot{e_{\rm i}}f_{\rm d,i}-\dot{e_{\rm o}}f_{\rm d,o}^{{}^{\prime}}=-(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A4)
where we assume $\sin{\varphi}\equiv\sin{\varphi_{\rm i}}=-\sin{\varphi_{\rm o}}$. The equation for $\ddot{\varphi}$ can be simplified to the form
$\ddot{\varphi}{=}-\omega_{0}^{2}\sin{\varphi},$ (A5)
where we take
$\omega_{0}^{2}=-3[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}).$ (A6)
The solution can be described as a libration of $\varphi$. The total energy $E$ of the system is the sum of the kinetic energy and potential energy. Hence $E=\frac{1}{2}\dot{\varphi}^{2}+2\omega_{0}^{2}\sin^{2}{\varphi/2}$. It is clear that the maximum libration occurs when $\dot{\varphi}=0$ at $\varphi=\pm\pi$. This implies $E_{\rm max}=2\omega_{0}^{2}$. We now set the value of $E$ equal to $E_{\rm max}$ and consider $\dot{\varphi}$ given by
$\dot{\varphi}=\pm 2\omega_{0}\cos{\varphi/2}.$ (A7)
We can relate the variation of $\dot{\varphi}$ to the variation in $e_{\rm i,o}$ by means of Eq. (A4). This gives
$d(e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{d,o^{{}^{\prime}}}){=}\pm\frac{(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})}{\omega_{0}}\sin{\frac{\varphi}{2}}d{\varphi}.$ (A8)
Substitute $\omega_{0}$ into Eq. (A8) and integrate it, we obtain
$e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}=-\frac{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}\cos{\frac{\varphi}{2}}.$ (A9)
The maximum eccentricity occurs when $\varphi=0$, which provides the eccentricity described in Sec. 2.1.1.
Appendix B: Extension of Stability and Escape Criteria to Arbitrary Mass Ratios
B.1: Escape Criterion
For the cases of $q\ll 1$ and $q\gg 1$, the system follows the same Hamiltonian equation:
$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (B1)
where
$\begin{cases}{\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}]^{2/3}\frac{e_{\rm i}^{2}}{4},\ \ \hfill[\mbox{$q\ll 1$}],\vspace{0.1cm}}\\ {\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm i}}\right)^{2/3}[\frac{j^{2}}{f_{\rm d,o}^{{}^{\prime}}}]^{2/3}\frac{e_{\rm o}^{2}}{4},\ \ \hfill[\mbox{$q\gg 1$}].}\end{cases}$ (B2)
Therefore, we hypothesize the existence of a form for $\Phi$ that satisfies the Hamiltonian equation and degenerates to Eq. (B3) in the extreme mass ratio cases:
$\Phi=\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}]^{2/3}\frac{e^{2}}{4},$ (B3)
where $e^{2}=e_{\rm i}^{2}+e_{\rm o}^{2}$. When $\eta_{\rm eq}{<}-3/2$ (equivalently $\Phi_{\rm eq}>1$), a separatrix point occurs and the system can cross this separatrix and escape from resonance (see Sec. 2.2.2). By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ (Eq. 4) into Eq. (B3), we obtain the general escape criterion as
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}.$ (B4)
where
$h(q)$ $=$ $\tau_{a}/\tau_{e}\left(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2}\right)$ (B5)
$=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$
B.2: Stability Criterion
The stability of the system decreases with the mass ratio $q$. When the outer planet is comparable or more massive than the inner planet ($q\lesssim 1$), the system may become overstable (see Eq. 22).However, all traps remain stable when the inner planet is more massive. Therefore, we introduce a factor $f(q)$ to fit the general stability criterion, which is given by
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q),$ (B6)
where
$f(q)=1-\frac{e_{\rm o,eq}^{2}}{e_{\rm i,eq}^{2}}\ =1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (B7)
When $q\ll 1$, $f(q)\sim 1$, the system is dominated by the outer planet. As $q$ increases, $f(q)$ decreases, indicating that the system becomes more stable. When $e_{\rm o,eq}>e_{\rm i,eq}$, $f(q)<0$, indicating that the system is dominated by the inner planet and all traps are stable traps.
B.3: Dependence on the mass ratio $q$
We run simulations to testify to our proposed overstability and escape criteria. Figure B.1 shows the numerical results of $\tau_{\rm a}/\tau_{\rm e}$ as a function of $q$. We fix $m_{\rm i}{=}10\ M_{\oplus}$ and set $\tau_{\rm a}{=}3\ (\tau_{\rm a})_{\rm slow,mig}$ to fulfill the trapping criteria. The solid lines denote the analytical criteria, while the colored dots represent the simulation results.
The numerical results agree pretty well with the analytical expectation at $q{\leq}1$. The simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes as $\tau_{\rm a}/\tau_{\rm e}$ decreases, and our analytical criteria separate these three regimes. In contrast, when $q\geq 1$, there are still two points in Figure B.1 where the simulations do not fully match the analysis.
First, the stability/overstability and escape criteria exhibit the opposite dependence $q$, with the two lines intersecting at $q{\sim}2$. For $q{\geq}2$, one would expect that the systems reside region below the overstability criterion (yellow line) correspond to escape. However, simulations show that all systems in this region are trapped overstably (no escape occurs). We attribute this to the fact that when $q{\geq}2$, the angular momentum deficit (AMD) (Laskar 1997) of the inner planet is larger than that of the outer planet, limiting the amplitude of oscillations in overstability (see Appendix B.4 for detailed derivation). As a result, the perturbations cannot grow sufficiently to cause escape from resonance. The dashed purple line in Figure B.1 represents the AMD criterion when the second-order terms of the eccentricities are considered. This new criterion is in good agreement with our numerical simulations.
Second, our overstable criterion suggests $\tau_{\rm a}/\tau_{\rm e}{<}0$ at $q{\geq}3$, indicating that all traps should be stable in this regime. However, in Figure B.1, it is evident that near $q\sim 4$ and $\tau_{a}/\tau_{e}\sim 10^{2}$, the planets remain in an overstability state. This is because we only considered the first-order term of the disturbing function. Second-order terms are neglected here, weakening the inner planet effect on the outer planet and leading to an overstable trap.
Figure B.1: Simulation results of $2:1$ MMR stability in $q-\tau_{a}/\tau_{e}$ parameter space with the mass of the inner planet $m_{\rm i}=10\ \times M_{\oplus}$. Green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate the planets eventually escape from the resonance. The lines represents our analytical criteria. It can be observed that our criteria overall align well with the numerical results.
B.4: Second-Order AMD Criterion
We expand the disturbing functions to the second order in $e_{\rm i}$ and $e_{\rm o}$, which gives
$R_{\rm D}^{\rm sec}$ $=$ $K_{1}(e_{\rm i}^{2}+e_{\rm o}^{2})+K_{2}e_{\rm i}e_{\rm o}\cos{(\varpi_{\rm i}-\varpi_{\rm o})},$ (B8)
$R_{\rm D}^{\rm res}$ $=$ $e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$ (B9)
$+[e_{\rm i}^{2}f_{3}\cos{\varphi_{3}}+e_{\rm i}e_{\rm o}f_{4}\cos{\varphi_{4}}+e_{\rm o}^{2}f_{5}\cos{\varphi_{5}}]\delta_{\rm j,2},$
$R_{\rm E}$ $=$ $-2e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2},$ (B10)
$R_{\rm I}$ $=$ $-\frac{1}{2}e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2},$ (B11)
where $\varphi_{3}{=}4\lambda_{\rm o}-2\lambda_{\rm i}-2\varpi_{\rm i}$, $\varphi_{4}{=}4\lambda_{\rm o}-2\lambda_{\rm i}-\varpi_{\rm i}-\varpi_{\rm o}$, $\varphi_{5}{=}4\lambda_{\rm o}-2\lambda_{\rm i}-2\varpi_{\rm o}$. Detailed expressions of $K_{1},K_{2},f_{3},f_{4}$ and $f_{5}$ can be found in Tables 8.1 and 8.4 of Murray & Dermott (1999). Lagrange’s equations can be written as
$\dot{\varpi_{\rm i}}$ $=$ $\frac{\mu_{\rm o}n_{\rm i}\alpha}{e_{\rm i}}[f_{\rm d,i}\cos{\varphi_{\rm i}}+(2e_{\rm i}f_{3}\cos{\varphi_{3}}+e_{\rm o}f_{4}\cos{\varphi_{4}})\delta_{\rm j,2}$ (B12)
$+2K_{1}e_{\rm i}+K_{2}e_{\rm o}\cos{(\varpi_{\rm i}-\varpi_{\rm o})}],$
$\dot{\varpi_{\rm o}}$ $=$ $\frac{\mu_{\rm i}n_{\rm o}}{e_{\rm o}}[f_{\rm d,o}^{{}^{\prime}}\cos{\varphi_{\rm o}}+(2e_{\rm o}f_{5}\cos{\varphi_{5}}+e_{\rm i}f_{4}\cos{\varphi_{4}})\delta_{\rm j,2}$ (B13)
$+2K_{1}e_{\rm o}+K_{2}e_{\rm i}\cos{(\varpi_{\rm i}-\varpi_{\rm o})}].$
The angular momentum deficit (AMD) (Laskar 1997) is given by
$AMD\sim\mu\sqrt{a}(1-\sqrt{1-e^{2}})\sim\mu\sqrt{a}e^{2}.$ (B14)
When the AMD of the inner planet exceeds that of the outer planet, the condition
$e_{\rm o}^{2}/e_{\rm i}^{2}>q\sqrt{\alpha},$ (B15)
is satisfied. Under this condition, the amplitude of oscillations following overstability is limited, and therefore the perturbations cannot grow sufficiently to cause escape. By substituting Eqs. (B12) and (B13) to eliminate either $e_{\rm i}$ and $e_{\rm o}$ under the condition $\dot{\varpi_{\rm i}}=\dot{\varpi_{\rm o}}$, we obtain the relationship between $e_{\rm i,o}$ and $q$. Assuming $e_{\rm i}$ takes the form: $e_{\rm i}^{2}={\tau_{e}}/{2j\tau_{a}}$, and substituting this into the derived form, we obtain our AMD criterion as
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm AMD}=\frac{2}{j}\left[\frac{(f_{3}-K_{1})+q^{{}^{\prime}2}(f_{5}+K_{1})}{q^{{}^{\prime}}f_{\rm d,o}^{{}^{\prime}}+f_{\rm d,i}}\right]^{2},$ (B16)
where $q^{{}^{\prime}2}\equiv q\sqrt{\alpha}$. The AMD criterion aligns well with the numerical results (Figure B.1).